# Continual Learning in Medical Image Analysis: A Comprehensive Review of Recent Advancements and Future Prospects

### v3

@@ 6.3 Episode performance and others
Zhang et al. (2023b) CT Abdomen multi- organ, tumor CIS Architectural Swin UNETR (Hatamizadeh et al. 2021) Supervised Average DSC BTCV (Landman et al. 2015), LiTS (Bilic et al. 2023), JHH (Xia et al. 2022) (private)
@@
Ji et al. (2023) 3D CT Whole-body organs CIS Architectural nnUNet Supervised DSC HD95 TotalSegmentator (Wasserthal et al. 2023), ChestOrgan, HNOrgan, EsoOrgan (3 private)
@@
Bera et al. (2023) MRI Prostate, hippocampus, spleen TIS, DIS Rehearsal Residual UNet Supervised DSC, average forgetting, BWT, accuracy (Prostate158 (Adams et al. 2022), NCI-ISBI (NCI ), Promise12 (Litjens et al. 2014), Decathlon (Antonelli et al. 2022)), (Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)), (Spleen dataset in Decathlon (NCI ))
@@
Zhu et al. (2023) CT, MRI Abdomen, muscles, prostate DIS Regularization U-Net Supervised DSC, ASSD Prostate (Liu et al. 2020; Bloch et al. 2015; Lemaître et al. 2015), abdomen (Mul ; Kavur et al. 2021), muscles (Zhu et al. 2021)
@@
Liu et al. (2023a) MRI Brain tumor CIS with domain shift Regularization ResNet-based 2D nnU-Net Supervised DSC, HD95 BraTS2013 (Menze et al. 2014), TCIA (Clark et al. 2013), CBICA (Bakas et al. 2018)
@@
Wang et al. (2023b) Endoscopic images Endoscopy, surgical instrument CIS Rehearsal + regularization ResNet101 Supervised mIoU EDD2020 (Ali et al. 2021, 2020), EndoVis18 (Allan et al. 2020), EndoVis17 (Allan et al. 2019)
Table 15: CL Literature for classification + segmentation
@@
Reference (year) Imaging modality Classification of CL scenario CL strategy Model Learning type Evaluation matrices Dataset
@@ 7 Discussion: Towards the future
Our exploration of the CL methods for medical imaging reveals a mix of challenges and strategies. The literature reflects a struggle with traditional paradigms, notably the need for explicit task/domain identification. While some methods demand prior knowledge of domain IDs, others explore task/domain-agnostic approaches that aim to eliminate this reliance Aljundi et al. (2017); Bayasi et al. (2021). The evolving notion of task/domain boundaries adds another layer of complexity, with real-world scenarios often defying the assumption of well-separated shifts. Non-rigid transitions and blurred boundaries, as explored by González et al. (2022); Wang et al. (2023a), challenge existing CL approaches. Sample annotation, a common requirement for CL in medical imaging, introduces resource challenges, especially in histopathology. While supervised strategies dominate, Ashfahani and Pratama (2022) showcase the feasibility of unsupervised CL, using minimal labeled samples. The evaluation of CL strategies involves a diverse set of metrics, including stability (backward transfer - BWT) and plasticity (forward transfer - FWT). These metrics, encompassing aspects like average forgetting and incremental learning, facilitate nuanced comparisons. In the classification domain, applications like chest X-rays and skin disease classification employ diverse CL strategies. Latent replay mitigates catastrophic forgetting but increases resource consumption, while knowledge distillation balances stability and plasticity Li et al. (2020); Srivastava et al. (2021). Segmentation tasks, spanning brain sclerosis lesion segmentation to cardiac segmentation, rely on regularization techniques like EWC and knowledge distillation Karthik et al. (2022); Li et al. (2022). The simultaneous handling of classification and segmentation tasks, evident in studies on dermoscopic image-based skin disease classification Li et al. (2020) and abdominal organ segmentation Liu et al. (2022), introduces a trade-off between resource efficiency and overall model performance. The synthesis of literature and comparative analyses unveils the current state of CL in medical imaging. However, challenges persist, and future research should explore unsupervised CL, address labeling costs, and refine evaluation metrics. As the field evolves, understanding the dynamic interplay between medical imaging tasks and CL methodologies will be crucial for driving innovation. In Tables 13, 14 and 15, we present the overview of the selected publications in the domain of classification, segmentation, and both combined, respectively. We highlight the domain/application, dataset information, imaging modalities, CL scenarios, and strategies, including the learning process and methodology adopted/proposed by different studies. Based on our critical analysis of existing literature, we discuss some of the open challenges and thus possible research directions for CL in the medical field.
@@
Ethical considerations and patient privacy: Ethical considerations and patient privacy concerns are paramount in CL for medical imaging. The use of patient data for model training raises ethical questions regarding consent, data anonymization, and potential biases. Future research, in alignment with ethical guidelines discussed by Chee et al. (2023), should prioritize the development of frameworks that ensure responsible data usage, privacy preservation, and ethical guidelines. Striking a balance between innovation and the protection of patient rights is imperative for the sustainable progress of CL in medical imaging.
@@
Continual learning in implicit neural representations: The exploration of CL within implicit neural representations Molaei et al. (2023) opens up new frontiers in leveraging the power of neural networks for dynamic learning scenarios. Implicit neural representations, commonly used in generative modeling and function approximation, can benefit immensely from CL techniques to adapt to new data and tasks seamlessly. As the data landscape evolves, implicit neural representations face challenges related to retaining past knowledge and efficiently incorporating new information Po et al. (2023); Yan et al. (2021b). CL offers a solution to these challenges by enabling models to update their representations while preserving previously acquired knowledge. This not only enhances the model’s ability to handle changing data distributions but also supports the development of intelligent systems that can learn and evolve over time. CL in implicit neural representations is instrumental for creating adaptable and intelligent models that can navigate the complexities of dynamic environments.

### v4

@@ 6.3 Episode performance and others
Zhang et al. (2023b) CT (a) Abdomen multi-organ segmentation, (b) abdomen to liver tumor segmentation CIS Architectural Swin UNETR (Hatamizadeh et al. 2021) Supervised Average DSC BTCV (Landman et al. 2015), LiTS (Bilic et al. 2023), JHH (Xia et al. 2022) (private)
@@
Ji et al. (2023) 3D CT Whole-body organ segmentation CIS Architectural nnUNet Supervised DSC HD95 TotalSegmentator (Wasserthal et al. 2023), ChestOrgan, HNOrgan, EsoOrgan (3 private)
@@
Bera et al. (2023) MRI Prostate segmentation, hippocampus segmentation, spleen segmentation TIS, DIS Rehearsal Residual UNet Supervised DSC, average forgetting, BWT, accuracy (Prostate158 (Adams et al. 2022), NCI-ISBI (NCI ), Promise12 (Litjens et al. 2014), Decathlon (Antonelli et al. 2022)), (Drayd (Denovellis et al. 2021), HarP (Boccardi et al. 2015)), (Spleen dataset in Decathlon (NCI ))
@@
Zhu et al. (2023) CT, MRI (a) Abdomen segmentation, (b) muscles segmentation, (c) prostate segmentation DIS Regularization U-Net Supervised DSC, ASSD Prostate (Liu et al. 2020a; Bloch et al. 2015; Lemaître et al. 2015), abdomen (Mul ; Kavur et al. 2021), muscles (Zhu et al. 2021)
@@
Liu et al. (2023a) MRI Brain tumor segmentation CIS + domain shift Regularization ResNet-based 2D nnU-Net Supervised DSC, HD95 BraTS2013 (Menze et al. 2014), TCIA (Clark et al. 2013), CBICA (Bakas et al. 2018)
@@
Wang et al. (2023b) Endoscopic images (a) Endoscopy segmentation, (b) surgical instrument segmentation CIS Rehearsal + regularization ResNet101 Supervised mIoU EDD2020 (Ali et al. 2021, 2020), EndoVis18 (Allan et al. 2020), EndoVis17 (Allan et al. 2019)
Table 20: Literature for CL based segmentation (Continued)
@@
Reference (year) Imaging modality Application CL scenario CL strategy Model Learning type Evaluation metric Dataset
Wei et al. (2023) MRI Brain tumor segmentation IIS Rehearsal UNet Supervised DSC LGG Segmentation (Buda et al. 2019)
Chen et al. (2023) Fundus, MRI (a) Optic disc segmentation, (b) cardiac segmentation, (c) prostate segmentation DIS Rehearsal + Regularization (KD) UNet Unsupervised DSC (a) REFUGE (Orlando et al. 2020), IDRiD (Porwal et al. 2018), RIM-ONE DL (Batista et al. 2020), (b) M&Ms challenge datasets (Campello et al. 2021), (c) RUNMC (Liu et al. 2020a), NCI-ISBI13 (Bloch et al. 2015), I2CVB (Lemaître et al. 2015), PROMISE12 (Litjens et al. 2014)
Li et al. (2024) MRI Cardiac segmentation DIS Rehearsal + Architecture Res-UNet Supervised DSC, BWT, FWT ACDC (Bernard et al. 2018), M&M (Campello et al. 2021)
Zhu et al. (2024b) MRI (a) Prostate segmentation (b) Cardiac segmentation DIS Regularization UNet Supervised HD95, DSC, BWT (a) RUNMC (Bloch et al. 2015), BMC (Bloch et al. 2015), I2CVB (Lemaître et al. 2015), UCL (Litjens et al. 2014), BIDMC (Litjens et al. 2014), HK (Litjens et al. 2014), (b) M&M (Campello et al. 2021)
Table 21: CL Literature for classification + segmentation and other application
Reference (year) Imaging modality Application CL scenario CL strategy Model Learning type Evaluation matrices Dataset
Zhang and Wang (2019b) Longitudinal MRI Alzheimer’s disease progression modelling (regression) TIS (7 Ep.) Regularization MLP Supervised wCC, PCC, rMSE ADNI-1 (Jack Jr et al. 2008),
@@
Wu et al. (2024) (PD, T1, T2) weighted MRI, X-ray Super-resolution TIS Regularization HAN (Niu et al. 2020) Supervised SSIM, PSNR, BWT, forgetting, Intransigence (plasticity) IXI (IXI ), Chest X-ray (Wang et al. 2017)
Ye et al. (2024) Multi-modality (medical report, MRI, X-ray, CT, histopathology) Representation learning TIS (5 Ep.) Rehearsal + Regularization Transformer Self-supervised DSC, HD95, AUC, accuracy, F1-score MIMIC-CXR 2.0.0 (Johnson et al. 2019), ADNI-1+ADNI-2+ADNI-GO (Jack Jr et al. 2008), DeepLesion (Yan et al. 2018), TCGA
Zhu et al. (2024a) Histopathology WSI retrieval CIS (4 Ep.) Rehearsal TransMIL (Shao et al. 2021) Supervised KRC, SRC, slide-level mean AP TCGA
@@ 7 Discussion: Towards the future
Our exploration of the CL methods for medical imaging reveals a mix of challenges and strategies. The literature reflects a struggle with traditional paradigms, notably the need for explicit task/domain identification. While some methods demand prior knowledge of domain IDs, others explore task/domain-agnostic approaches that aim to eliminate this reliance Aljundi et al. (2017); Bayasi et al. (2021). The evolving notion of task/domain boundaries adds another layer of complexity, with real-world scenarios often defying the assumption of well-separated shifts. Non-rigid transitions and blurred boundaries, as explored by González et al. (2022); Wang et al. (2023a), pose challenges to existing CL approaches. Sample annotation, a common requirement for CL in medical imaging, introduces resource challenges, especially in histopathology. While supervised strategies dominate, Ashfahani and Pratama (2022) showcase the feasibility of unsupervised CL, using minimal labeled samples. The evaluation of CL strategies involves a diverse set of metrics, including stability (backward transfer - BWT) and plasticity (forward transfer - FWT). These metrics, encompassing aspects like average forgetting and incremental learning, facilitate nuanced comparisons. In the classification domain, applications like chest X-rays and skin disease classification employ diverse CL strategies. Latent replay mitigates catastrophic forgetting but increases resource consumption, while knowledge distillation balances stability and plasticity Li et al. (2020a); Srivastava et al. (2021). Segmentation tasks, spanning brain sclerosis lesion segmentation to cardiac segmentation, rely on regularization techniques like EWC and knowledge distillation Karthik et al. (2022); Li et al. (2022). The simultaneous handling of classification and segmentation tasks, evident in studies on dermoscopic image-based skin disease classification Li et al. (2020a) and abdominal organ segmentation Liu et al. (2022), introduces a trade-off between resource efficiency and overall model performance. The synthesis of literature and comparative analyses unveils the current state of CL in medical imaging. However, challenges persist, and future research should explore unsupervised CL, address labeling costs, and refine evaluation metrics. As the field evolves, understanding the dynamic interplay between medical imaging tasks and CL methodologies will be crucial for driving innovation. In Tables 17, 19 and 21, we present the overview of the selected publications in the domain of classification, segmentation, and both combined, respectively. We highlight the domain/application, dataset information, imaging modalities, CL scenarios, and strategies, including the learning process and methodology adopted/proposed by different studies. From the last year (2023), we also witness some exploration in non-imaging medical applications including disease classification, gesture classification, arrhythmia detection, etc. Further, apart from classification and segmentation, applications like super-resolution Wu et al. (2024), WSI retrieval Zhu et al. (2024a), representation learning Ye et al. (2024), etc., also benefit from CL. Based on our critical analysis of existing literature, we discuss some of the open challenges and thus possible research directions for CL in the medical field.
Non-imaging or other applications Recently, there have also been explorations in wearable medical sensor data. For example, Hua et al. (2023) explore experience replay (various sample selection strategies were explored) and knowledge distillation (on old data) based CL for incrementally learning gesture classes with sEMG data. Similarly, Sun et al. (2023b) and Sun et al. (2023c) incrementally learn classes from time series signals (ECG, categorical data) for applications including time-series health monitoring, mortality identification, sepsis identification, gesture classification, etc. Further, there have been studies to handle domain shifts due to multi-site data Kim et al. (2024), patient-level splits Li and Jha (2023), data distributional shifts (change in mean and standard deviation) Aslam et al. (2024) in ECG, physiological signal, or other time-series data for disease classification and arrhythmia detection.
Largely CL has been explored for mainstream applications such as classification and segmentation, but some other application, such as image reconstruction, registration, translation, generation, anomaly detection, etc., can also benefit from CL techniques. Recently, Wu et al. (2024) explored CL for the image super-resolution task across different imaging modalities including PD, T1, and T2 weighted MRI, and X-ray is learned in a continual manner with a single super-resolution model. Notably, to deal with forgetting, authors use the idea of constraining gradient updates according to the importance of parameters as well as adding a distillation loss as a regularization.
@@
Ethical considerations and patient privacy: Ethical considerations and patient privacy concerns are paramount in CL for medical imaging. The use of patient data for model training raises ethical questions regarding consent, data anonymization, and potential biases. Future research, in alignment with ethical guidelines discussed by Chee et al. (2023), should prioritize the development of frameworks that ensure responsible data usage, privacy preservation, and ethical guidelines. Striking a balance between innovation and the protection of patient rights is imperative for the sustainable progress of CL in medical imaging. Verma et al. (2023) stress over exemplar-free CL approaches for colon pathology and retina data-based disease classification. They present a comparative study of 12 existing approaches in all three major categories offering privacy-aware solutions. Specifically, they use LwF Li and Hoiem (2018), LwM Dhar et al. (2019), EWC Kirkpatrick et al. (2017), SI Zenke et al. (2017), MAS Aljundi et al. (2017), MUC-MAS Liu et al. (2020b), RWalk Chaudhry et al. (2018b), OWM Zeng et al. (2019), and GPM Saha et al. (2021) in regularization category, GR Shin et al. (2017) and BIR Van de Ven et al. (2020) in generative replay category, and EFT Verma et al. (2021) in architecture-based category.
Limitations with Gigapixel WSIs: The high resolution of whole slide images (WSIs), often 50,000 x 50,000 pixels, presents significant computational challenges for deep learning model design. The variability in WSI imaging technology and staining protocols affects model performance on new data, requiring continual adaptation of WSI analysis methods Lai et al. (2024). Directly applying standard CL approaches to hierarchical WSI models can lead to severe knowledge forgetting of previously seen datasets. Additionally, WSIs are gigapixel images with only slide-level labels, making storage and computation for rehearsal-based CL impractical due to limited memory. Huang et al. (2023) proposed a CL method, "ConSlide", to progressively update a hierarchical WSI analysis architecture using sequentially acquired heterogeneous WSI datasets. In this approach, a representative set of past datasets is stored and periodically reorganized and replayed during model updates using an asynchronous updating mechanism. Thandiackal et al. (2024) proposed a rehearsal-based domain incremental scenario with unsupervised learning for tissue classification in WSIs patches. They utilize the generative feature-driven image replay in conjunction with a dual-purpose discriminator. Further, Zhu et al. (2024a) enhance reservoir sampling for WSI retrieval by incorporating distance consistency-based rehearsal.
@@
Continual learning in implicit neural representations: The exploration of CL within implicit neural representations Molaei et al. (2023) opens up new frontiers in leveraging the power of neural networks for dynamic learning scenarios. Implicit neural representations refer to a class of models where functions, such as geometric shapes or scenes, are represented as continuous signals rather than discrete entities Sitzmann et al. (2020). These representations are commonly used in generative modeling and function approximation due to their ability to capture complex, high-dimensional data in a compact form Molaei et al. (2023). Implicit neural representations can benefit immensely from CL techniques to adapt to new data and tasks seamlessly. As the data landscape evolves, implicit neural representations face challenges related to retaining past knowledge and efficiently incorporating new information Po et al. (2023); Yan et al. (2021b). CL offers a solution to these challenges by enabling models to update their representations while preserving previously acquired knowledge. This not only enhances the model’s ability to handle changing data distributions but also supports the development of intelligent systems that can learn and evolve over time. CL in implicit neural representations is instrumental in creating adaptable and intelligent models that can navigate the complexities of dynamic environments.